OpenStudy (anonymous):

Find the sum of each finite geometric series. 1+2+4+...+2048

OpenStudy (kropot72):

The common ratio r is 2 in this case. When the common ratio is positive and greater than 1 the formula for the sum of n terms is \[S _{n}=\frac{a(r ^{n}-1)}{r-1}\] For this question we need to find the value of n for the last term 2048. The nth term is found from the formula \[n ^{th} term=ar ^{n-1}\] Plugging in the known values of a and r gives \[2048=1\times2^{n-1}\ .........(1)\] 2 to the power of 11 = 2048. Therefore n = 11 - 1 = 10 Now you just need to plug the values of a, r and n into the formula for the sum of n terms.

OpenStudy (anonymous):

the last part is the part I don't understand, can you explain it simply what I have to do "to plug in"

OpenStudy (kropot72):

Sorry, my bad. The final part of my post should be: "2 to the power of 11 = 2048. Therefore n = 11 + 1 = 12. Now you just need to plug the values of a, r and n into the formula for the sum of n terms." The values are as follows: a = 1 r = 2 n = 12 The formula is \[S _{n}=\frac{a(r ^{n}-1)}{r-1}\]

OpenStudy (anonymous):

I thought it was n-1 so wouldn't it be 11 - 1 instead of 11 + 1

OpenStudy (kropot72):

\[2^{11}=2048\] \[2^{n-1}=2048\] Therefore n - 1 = 11 ........(2) Adding 1 to both sides of equation (2) gives n = 11 + 1 = 12

OpenStudy (anonymous):

oh okay!

OpenStudy (anonymous):

can you work me through the formula?

OpenStudy (kropot72):

The values are as follows: a = 1 r = 2 n = 12 The formula is \[S _{n}=\frac{a(r ^{n}-1)}{r-1}\] So after plugging in the values we get \[S _{12}=\frac{1(2^{12}-1)}{2-1}=you\ can\ calculate\]

OpenStudy (anonymous):

that's actually the part I don't get, don't understand how to calculate it :(

OpenStudy (anonymous):

2^18 = 262144?

OpenStudy (kropot72):

Where did you get '2 to the power of 18' from?

OpenStudy (anonymous):

oh gosh, I misread it.

OpenStudy (anonymous):

oh gosh, I misread it.

OpenStudy (anonymous):

the 12 looked like an 18 to me, I'm sorry, haha

OpenStudy (anonymous):

so 4096

OpenStudy (kropot72):

When you simplify the fraction you get \[S _{12}=\frac{1(2^{12}-1)}{2-1}=2^{12}-1\]

OpenStudy (anonymous):

4095 is it then?

OpenStudy (kropot72):

4095 is correct!

OpenStudy (anonymous):

thank you so much! I actually think I'm starting to get it, I just need nudges in the right direction! :)

OpenStudy (kropot72):

You're welcome :) Glad to hear you are understanding.